elaborate on the meaning of law of increasing opportunity cost. i have read the definition many times but need a more “normal” explanation!

give realistic or practical examples for the concept.

thank you
Answer #1

The law shows that with each additional unit being produced has increasingly higher opportunity cost. Here, the opportunity cost is the quantity of output foregone of one type of output ( example burger) , if one unit of additional output of another good ( example coffee)  is produced. It can be further understood by the following example.

An economy has following production possibility combination:

S.No. Coffee Burger
1 0 10
2 1 9
3 2 7
4 3 4
5 4 0

An economy only produces two goods that are coffee and burger as shows above.

At point 1, the economy is producing no coffee and 10 units of burger. When economy produces 1 unit of coffee, then output of burger decreases to 9. It means that opportunity cost of first unit of coffee is 1 unit of burger (10-9 = 1).

For second unit of coffee produced, the burger output decreases to 7.

Now, for second unit of coffee, the opportunity cost = 9-7 = 2.

For third unit of coffee produced, the burger output decreases to 4.

Now, for third unit of coffee, the opportunity cost = 7-4 = 3.

So, it can be seen that opportunity cost is growing from 0 to 1, then 1 to 2 and then 2 to 3, with each additional unit of coffee production. It is due to the law of increasing opportunity cost.